# Enhancing the Performance of Aspect-Based Sentiment Analysis Systems

### v2

@@ 3.3 Transformer Network
where ${\bf Q}$, ${\bf K}$ and ${\bf V}$ represent the query, key and value matrices with dimensions $d_{k}$, $d_{k}$ and $d_{v}$, respectively. Usually, a transformer employs a multi-head attention mechanism to project the query, key and value matrices $h$ times to provide improved internal contextual associations. The mathematical expression for this process is as follows:
@@
where ${\bf W,W}_{Q},{\bf W}_{K}$, and ${\bf W}_{V}$ indicate weight matrices of the self-attention mechanism. Next, the transformer encoder applies two normalization layers to extract deep features of the context. Consider ${\bf Z}_{out}$ as the result matrix generated by the transformer encoder, and its formulation can be expressed as follows:
@@ 3.4 Dependency-parsing Layer
The output hidden vector of the Bi-LSTM network, denoted as ${\bf h}_{t}$, can be used as an input to the dependency-parsing layer in the Bi-GCN, which helps to identify word dependencies in a sentence by constructing a dependency tree. Formally, the Bi-GCN initiates a matrix to record the edge information weights. Initially, it derives the adjacency matrix ${\bf A}\in\mathbb{R}^{n\times n}$ from the sentence’s dependency tree. The conventional form of this adjacency matrix can be expressed as follows:
${\bf A}_{ij}=\begin{cases}1,\quad i=j.\\ 1,\quad i\neq j,~{}i~{}\text{and}~{}j~{}\text{have dependencies}.\\ 0,\quad\text{otherwise}.\end{cases}$ (11)
GCN lacks the ability to capture rich edge information due to the discrete nature of the adjacency matrix ( comprising only 0s and 1s). We present an adjacency matrix that contains non-discrete edge information, which is formulated as follows:
@@
Here, SDI is used to compute syntactic dependency information between the word $i$ and the word $j$. The computation is summarized below:
$\text{SDI}(i,j)=\frac{\text{Count}(\text{sd}(i,j))}{\text{Count}(\text{sd}(\cdot))},$ (13)
where $\text{sd}(i,j)$ and $\text{sd}(\cdot)$ represent the syntactic relationship between word pair $i$ and $j$ and the overall syntax of the dataset.
@@ 3.5 Bi-GCN Network
A GCN is a specialized variant of a Convolutional Neural Network (CNN) specifically tailored for encoding structured graph data. In the scenario of a textual graph consisting of $n$ words, we generate an adjacency matrix based on the syntactic matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ to capture information regarding syntactic dependencies. The representation of word $i$ at layer $l$ is symbolized as ${\bf h}^{l}_{i}$. This graphical structure is conceptually depicted as follows:
@@
In the GCN, $\mathbf{W}^{l}$ serves as the weight for linear transformation, $\mathbf{b}^{l}$ acts as the bias, and $\sigma$ represents the nonlinear activation function. The GCN harnesses the sentence’s dependency tree to enforce syntactic constraints, facilitating the identification of descriptive words associated with a specific aspect, taking into account their syntactic proximity. This capability equips GCN to effectively address situations where words related to the sentiment of an aspect are not adjacent. As a result, GCN is a well-suited choice for conducting aspect-based sentiment analysis. The update of each word’s representation is executed through the utilization of Bi-GCN in as follows:
$\overrightarrow{\bf h}^{l}_{i}=\sum_{j=1}^{n}\tilde{\mathbf{A}}_{ij}\mathbf{W}^{l}h_{j}^{l-1},$ (15)
$\overleftarrow{\bf h}^{l}_{i}=\sum_{j=1}^{n}{\tilde{\mathbf{A}}_{ij}}^{T}\mathbf{W}^{l}h_{j}^{l-1},$ (16)
where $\overrightarrow{\bf h}^{l}_{i}$ and $\overleftarrow{\bf h}^{l}_{i}$ indicate the outputs of the forward and backward hidden layers for the word $i$ in the $l$th layer. Then, the forward and backward representations can be concatenated by
$\tilde{\bf h}^{l}_{i}=\overrightarrow{\bf h}^{l}_{i}\oplus\overleftarrow{\bf h}^{l}_{i},$ (17)
and
${\bf h}^{l}_{i}=\text{ReLU}(\frac{\tilde{\bf h}^{l}_{i}}{d_{i}+1}\mathbf{W}^{l}+\mathbf{b}^{l}),$ (18)
@@ 3.6 Aspect-Specific Masking
We utilize the masking technique to reduce redundancy in the Bi-GCN hidden representation by hiding aspect-independent text. To be more specific, this masking can be depicted as follows:
${\bf h}_{i}^{l}=0,~{}1\leq i<\gamma,~{}\gamma+m-1<t\leq n.$ (19)
The result from the zero mask layer consists of aspect-focused characteristics.
$\mathbf{H}_{mask}^{l}=\left\{{\bf 0},\cdots,{\bf h}_{\gamma}^{l},\cdots,{\bf h}_{\gamma+m-1}^{l},\cdots,{\bf 0}\right\}.$ (20)
By employing graph convolution, the feature vector $\mathbf{H}_{mask}^{l}$ adeptly captures contextual information surrounding specific aspects. This encompassing process considers both syntactic dependencies and extended multi-word relationships. Moreover, an attention mechanism is harnessed to produce a vector representation of the concealed layer. The calculation of attention weights is executed as follows:
$\beta_{i}=\sum_{j=1}^{n}{\bf h}_{j}^{T}{\bf h}_{j}^{l},$ (21)
$\alpha_{i}=\frac{\exp(\beta_{i})}{\sum_{j=1}^{n}\exp(\beta_{j}))}.$ (22)
Finally, the sentiment classification process combines the cumulative representation scores mentioned earlier with the attention scores generated by the transformer. This results in the following sentiment classification:
@@ 3.7 Sentiment Classification
The $\mathbf{r}_{out}$ computed in the previous step is used as input to the fully connected layer for sentiment classification. The formula is as follows:
$\mathbf{prob}=\text{softmax}(\mathbf{res}_{out}\mathbf{W}_{p}+\mathbf{b}).$ (24)
3.8 Training Phase
Utilizing both cross-entropy and $\mathit{L_{2}}$ regularization [46], the loss function for SentiSys is formulated as follows:
@@ 3.8 Training Phase
where $\mathit{p}$ signifies the label, and $C$ represents the dataset. $\mathbf{prob_{\hat{\mathit{p}}}}$ denotes the $\hat{\mathit{p}}$th element of $\mathbf{prob}$. $\Theta$ and $\lambda$ are the trainable parameters and the coefficient for $\mathit{L_{2}}$ regularization, respectively. To provide a clearer insight into the SentiSys model, Algorithm 1 illustrates the training and testing procedures of the model, facilitating a more comprehensive understanding.
// Read sample sentences from datasets.
1 Tokenize a sentence with Eq.(1).
2 Embed words in word embedding layer.
3 Analyze dependency information of the sentence.
// A Bi-LSTM layer.
4 Use a Bi-LSTM layer to capture the contexual features of the sample sentence with Eqs. (2),(3), and (4).
// A transformer layer.
5 Encode positions with Eqs. (5) and (6).
6 Employ a transformer encoder architecture to capture the global features within a sentence with Eqs. (7), (8), (9), and (10).
// Dependency-parsing layer.
7 Replace the weight matrix (11) with Eq. (12) and save more available edge information for the graph.
Compute SDI with Eq. (13). // A Bi-GCN submodule.
8 Use Bi-GCN to extract the features of the structure graph with Eqs. (14), (15), (16), (17), and (18).
// Aspect-specific masking.
9 Utilize aspect-specific masking techniques to reduce redundant information in the Bi-GCN with Eqs. (19), (20), (21), (22), and (23).
// Sentiment classification.
10 Apply fully connected layers to construct a classifier to categorize the sentiment with Eq. (24).
// Training loss.
11 Train SentiSys with Eq. (25).
// Testing phase.
Evaluate the performance of SentiSys on test datasets.
Algorithm 1 Algorithm for the proposed SentiSys.
@@ 4 Experiments
The efficacy of the SentiSys model is substantiated through experiments on four widely employed benchmark datasets. Furthermore, the influence of GCN layers on performance is deliberated. Lastly, ablation studies are conducted to examine syntactic dependency, the edge weight matrix (SDI), and bi-direction.
@@ 4.1 Datasets
Four commonly used benchmark datasets were employed to assess the performance of the SentiSys model. The Twitter dataset [47] encompasses 6,940 Twitter posts from users, with 6,248 data points for training and 692 for testing. Additionally, three datasets, namely Rest14 [48], Rest15 [49], and Rest16 [50] contain customer comments and restaurant rating information. Rest14 comprises 4,728 customer comments, with 3,608 for training and 1,120 for testing. Rest15 includes 1,746 reviews, with 1,204 for training and 542 for testing. Rest16, the 2016 benchmark dataset, consists of 2,364 reviews, of which 1,748 are used for training, and the remaining 616 are for testing. To enhance dataset quality, sentences with conflicting polarity or unclear meaning were removed from the Rest15 and Rest16 datasets. A comprehensive overview of the four datasets’ statistics is presented in Table 1.
Table 1: Comprehensive overview of the four datasets’ statistics.
@@ 4.2 Experimental Setup
In the experiments, we initialize the word embedding layer using a pre-trained GloVe vectors [51] with a dimensionality of 300. The hidden layer of the Bi-LSTM also has a dimensionality of 300. The Bi-GCN of SentiSys contains 3 layers with parameters initialized using a uniform distribution. The optimizer is Adam [52], which has a learning rate of 0.001. We use $\mathit{L_{2}}$ regularization with a batch size of 32. The training process is up to 100 epochs.
@@
4.3 Baselines for Comparison
We first evaluate the performance of SentiSys by comparing it to baseline models. These baseline models are described below:
  • SVM: The machine learning-based support vector machine (SVM) is widely used in SemEval 2014 Task 4. This model utilizes a conventional methodology for the extraction of essential features, making it a well-established approach in natural language processing and sentiment analysis tasks.
  • LSTM [35]: This methodology leverages the LSTM’s ability to capture and encode sequential information, with the focus on the hidden layer vector at the last time step, which is instrumental in making sentiment predictions.
  • TD-LSTM [35]: This specialized LSTM variant is designed to extract interaction features that capture the intricate relationship between the target sentiment, often a specific aspect or entity, and the entire contextual information.
  • MemNet [53]: MemNet improves the quality of sentence representations by efficiently integrating external textual information.
  • AOA [54]: The attention-over-attention (AOA) mechanism-based architecture plays a pivotal role in enhancing the classification process. AOA’s unique approach involves employing attention mechanisms not just once but in a layered fashion, enabling more refined and context-aware classification.
  • IAN [8]: The interactive attention networks (IAN) present a novel approach to information extraction. IAN employs a dual LSTM architecture to independently capture representations for aspect terms and contextual information.
  • TNet-LF [36]: The transformation network with local feedback (TNet-LF) introduces a forward context-preserving transformation mechanism. This innovative approach serves a dual purpose: to retain and amplify the valuable information within the context. TNet-LF achieves this by implementing a transformation process that not only safeguards but also enriches the contextual information, ultimately enhancing the overall understanding of the data.
  • ASGCN [40]: The pioneering aspect-specific graph convolutional network (ASGCN) marks a significant milestone in the field of sentiment analysis by harnessing the power of graph convolution, a novel approach to analyzing and understanding sentiment with a focus on specific aspects. By utilizing the capabilities of a GCN, ASGCN enhances the accuracy and effectiveness of sentiment analysis in the context of different aspects, making it a noteworthy advancement in this domain.
  • MGAN [9]: The Multi-grained attention network (MGAN) was introduced as an innovative approach with the primary objective of discerning and capturing the nuanced sentiments expressed in a given text, particularly in relation to specific aspects or entities.
  • Sentic LSTM [37]: Sentic LSTM incorporates a trio of key components to enhance its sentiment analysis capabilities. These components include an aspect-specific attention mechanism, a module for embedding commonsense knowledge, and an attentive LSTM.
  • CMLA-ALSTM [55]: This network incorporates multiple layers of attention mechanisms, with tensor operators implemented within each layer. The use of tensor operators at each layer enhances the network’s ability to capture and process complex relationships and patterns within the data. This design enables the network to perform intricate computations and feature extraction, making it a powerful tool for tasks that require advanced attention mechanisms and data analysis.
  • RACL [38]: Relation-aware collaborative learning (RACL) is a methodology that enables subtasks to synchronize their efforts within a network through the process of relational propagation. RACL’s fundamental concept is the establishment of connections among various subtasks, enabling them to effectively communicate and collaborate.
  • AEGCN [41]: Attention-based encoding GCN (AEGCN) leverages a multi-headed attention mechanism in conjunction with an enhanced GCN for sentiment analysis. This sophisticated approach employs multiple attention heads to focus on various aspects of the input data and combines them with an improved GCN, resulting in a more powerful method for sentiment analysis.
  • SK-GCN [42]: The syntax- and knowledge-based GCN (SK-GCN) harnesses the structure of syntactic dependency trees and the wealth of commonsense knowledge by integrating them with a GCN.
  • AGCN [43]: The aggregated GCN (AGCN) effectively utilizes node feature information through a process of iteratively enhancing the representation of individual nodes based on their local neighborhoods.
@@ 4.4 Evaluation Measures
We use two performance metrics, Accuracy (Acc) and Macro F1 (F1), to evaluate and validate the effectiveness of the proposed SentiSys. Acc and F1 are widely recognized in the field of multi-classification tasks to assess the predictive and classification capabilities of models. Concretely, Acc quantifies the ratio of correctly classified samples and thus directly measures the accuracy of the model. F1 takes into account both accuracy and recall, offering a comprehensive perspective on the model’s performance. Formally, these two evaluation measures are calculated as follows:
$\text{Acc}=\frac{TP+TN}{TP+TN+FP+FN},$
$\text{Precision}=\frac{TP}{TP+FP},$
$\text{Recall}=\frac{TP}{TP+FN},$
[…]
where TP, TN, FP, and FN refer to the correctly identified positive cases, correctly identified negative cases, incorrectly identified positive cases, and incorrectly identified negative cases.
@@ 4.5 Evaluation Results
Table 2 presents a comparative analysis of the Acc and F1 scores between the proposed SentiSys model and 15 other baseline models. Notably, the SentiSys model demonstrates superior performance across various metrics in the Twitter, Rest14, and Rest15 datasets. While the AGCN model achieved the highest Acc and F1 scores in the Twitter dataset, it did not exhibit the same level of effectiveness as the SentiSys model in the Rest14 and Rest16 datasets. Specifically, on the Rest14 dataset, the SentiSys model stands out, achieving significantly higher Acc and F1 scores compared to the other baseline models. In the Rest15 dataset, RACL achieved the highest Acc score, but its F1 score was relatively low at 59.85. When it comes to the Rest16 dataset, the Acc and F1 scores of TNet-LF were only marginally superior to those of the SentiSys model. This observation could be attributed to the critical role of information from both parent and child nodes, highlighting the significance of considering the entire graph structure rather than solely processing it as a directed graph. Additionally, the limited sentence grammar in the Twitter dataset may have constrained prediction results. In summary, the experimental results show that our proposed SentiSys enhances the performance of aspect-based sentiment analysis systems compared with other baseline models, especially for TNet-LF on the Rest16 dataset. Furthermore, SentiSys outperforms ASGCN on all four benchmark datasets, which highlights the impact of word dependence on enhancing the classification results.
@@ 4.6 Effect of GCN Layers
Figures 3 and 4 show the effects of different GCN layers on the Acc and F1 scores of the proposed SentiSys model on the Rest14 dataset. X and Y axes indicate the number of GCN layers and the corresponding Acc and F1 scores, respectively. The experimental results show that the Acc and F1 scores tended to increase and then decrease as the number of layers increased. The Acc and F1 scores achieved the maximum values with the number of GCN layers of 3. The scores are 0.818 and 0.731, respectively.
4.7 Ablation Studies
Table 3: Ablation studies on syntactical dependencies ($\boldsymbol{D}$), edge weights ($\boldsymbol{EW}$), and bidirectional ($\boldsymbol{Bi}$) structures for the SentiSys.
Dataset Measures $\boldsymbol{D}$ w/o $\boldsymbol{EW}$ w/o $\boldsymbol{Bi}$ w/o SentiSys
Twitter Acc (%) $\uparrow$ 71.53 71.39 71.80 72.40
F1 (%) $\uparrow$ 69.68 69.37 70.42 70.67
Rest14 Acc (%) $\uparrow$ 80.86 81.34 81.57 81.70
F1 (%) $\uparrow$ 72.19 71.70 73.35 73.63
Rest15 Acc (%) $\uparrow$ 79.34 79.15 80.44 79.34
F1 (%) $\uparrow$ 60.78 60.06 61.58 61.99
Rest16 Acc (%) $\uparrow$ 88.69 88.28 88.37 88.80
F1 (%) $\uparrow$ 66.64 67.03 69.20 68.96
    The values in gray cells indicate the maximum values.
The three critical parts of the proposed SentiSys model are the grammar dependency tree, the edge weight matrix, and the Bi-GCN. Ablation studies are conducted to verify the effects of each submodule on performance. Table 3 shows the Acc and F1 scores of the proposed SentiSys model without grammar dependency, with the discrete edge weight matrix, and without the bidirectional GCN on the four benchmark datasets. The results show that the grammar dependency technique, which leads to a significant improvement in the model performance. The SentiSys model with dependencies outperformed the results without dependencies on all four datasets. Similarly, the results with the SDI edge weight matrix are better than the traditional edge weight matrix with the discrete number representations (i.e., 0 and 1) on all four datasets. It indicates that the edge weight matrix with SDI improves the performance. Additionally, the SentiSys model with the addition of Bi-GCN outperformed the results without bi-direction on six of the eight measures. Overall, the syntactic dependency tree, the edge weight matrix, and Bi-GCN effectively improved the model performance.

### v3

@@ 3.3 Transformer Network
where ${\bf Q}$, ${\bf K}$ and ${\bf V}$ represent the query, key and value matrices with dimensions $d_{k}$, $d_{k}$ and $d_{v}$. ${\bf W,W}_{Q},{\bf W}_{K}$, and ${\bf W}_{V}$ indicate weight matrices of the self-attention mechanism. Next, the transformer encoder applies two normalization layers to extract deep features of the context. Consider ${\bf Z}_{out}$ as the result matrix generated by the transformer encoder, and its formulation can be expressed as follows:
@@ 3.4 Dependency-parsing Layer
The output hidden vector of the Bi-LSTM network, denoted as ${\bf h}_{t}$, can be used as an input to the dependency-parsing layer in the Bi-GCN, which helps to identify word dependencies in a sentence by constructing a dependency tree. Formally, the Bi-GCN initiates a matrix to record the edge information weights. Initially, it derives the adjacency matrix ${\bf A}\in\mathbb{R}^{n\times n}$ from the sentence’s dependency tree. We present an adjacency matrix that contains non-discrete edge information, formulated as
@@
Here, SDI is used to compute syntactic dependency information between the word $i$ and the word $j$. The computation is calculated as $\text{SDI}(i,j)=\frac{\text{Count}(\text{sd}(i,j))}{\text{Count}(\text{sd}(\cdot))}$, where $\text{sd}(i,j)$ and $\text{sd}(\cdot)$ represent the syntactic relationship between word pair $i$ and $j$ and the overall syntax of the dataset.
@@ 3.5 Bi-GCN Network
A GCN is a specialized variant of a Convolutional Neural Network (CNN) specifically tailored for encoding structured graph data. In the scenario of a textual graph consisting of $n$ words, we generate an adjacency matrix based on the syntactic matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ to capture information regarding syntactic dependencies. Word $i$ at layer $l$ can be represented as ${\bf h}^{l}_{i}$. This graphical structure is depicted as
@@
In GCN, $\mathbf{W}^{l}$ is the weight for linear transformation, $\mathbf{b}^{l}$ acts as bias, and $\sigma$ is the nonlinear activation function. GCN utilizes the sentence’s dependency tree to enforce syntactic constraints, facilitating the identification of descriptive words associated with a specific aspect, considering their syntactic proximity. This capability enables GCN to effectively handle non-adjacent words related to aspect sentiment, making it a suitable choice for aspect-based sentiment analysis. The update of each word’s representation is executed through Bi-GCN as follows:
$\overrightarrow{\bf h}^{l}_{i}=\sum_{j=1}^{n}\tilde{\mathbf{A}}_{ij}\mathbf{W}^{l}h_{j}^{l-1}~{}~{}\text{and}~{}~{}\overleftarrow{\bf h}^{l}_{i}=\sum_{j=1}^{n}{\tilde{\mathbf{A}}_{ij}}^{T}\mathbf{W}^{l}h_{j}^{l-1},$ (12)
where $\overrightarrow{\bf h}^{l}_{i}$ and $\overleftarrow{\bf h}^{l}_{i}$ are the outputs of the forward and backward hidden layers for the word $i$ in the $l$th layer. The forward and backward representations are concatenated by
$\tilde{\bf h}^{l}_{i}=\overrightarrow{\bf h}^{l}_{i}\oplus\overleftarrow{\bf h}^{l}_{i}~{}~{}\text{and}~{}~{}{\bf h}^{l}_{i}=\text{ReLU}(\frac{\tilde{\bf h}^{l}_{i}}{d_{i}+1}\mathbf{W}^{l}+\mathbf{b}^{l}),$ (13)
@@ 3.6 Aspect-Specific Masking
We utilize the masking technique to reduce redundancy in the Bi-GCN hidden representation by hiding aspect-independent text. This masking can be depicted as ${\bf h}_{i}^{l}=0,~{}1\leq i<\gamma,~{}\gamma+m-1<t\leq n$. The result from the zero mask layer consists of aspect-focused characteristics, represented as $\mathbf{H}_{mask}^{l}=\left\{{\bf 0},\cdots,{\bf h}_{\gamma}^{l},\cdots,{\bf h}_{\gamma+m-1}^{l},\cdots,{\bf 0}\right\}$. Utilizing graph convolution, $\mathbf{H}_{mask}^{l}$ adeptly captures contextual information, including syntactic dependencies and extended multi-word relationships. The attention mechanism for the hidden layer operates as follows:
$\beta_{i}=\sum_{j=1}^{n}{\bf h}_{j}^{T}{\bf h}_{j}^{l}~{}~{}\text{and}~{}~{}\alpha_{i}=\frac{\exp(\beta_{i})}{\sum_{j=1}^{n}\exp(\beta_{j}))}.$ (14)
Finally, the sentiment classification combines cumulative representation scores with attentions from the transformer, resulting in the sentiment classification:
@@ 3.7 Sentiment Classification
The computed $\mathbf{r}{out}$ from the previous step serves as input to the fully connected layer for sentiment classification, given by $\mathbf{prob}=\text{softmax}(\mathbf{r}{out}\mathbf{W}_{p}+\mathbf{b})$. Finally, utilizing cross-entropy and $\mathit{L_{2}}$ regularization [7], the loss function is as follows:
@@
where $\mathit{p}$ signifies the label, and $C$ represents the dataset. $\mathbf{prob_{\hat{\mathit{p}}}}$ denotes the $\hat{\mathit{p}}$th element of $\mathbf{prob}$. $\Theta$ and $\lambda$ are the trainable parameters.
@@ 4.1 Datasets
Table 1: Statistics of the four datasets.
@@
Four commonly used benchmark datasets were employed to assess the performance of the SentiSys model. The Twitter dataset [10] encompasses 6,940 Twitter posts from users, with 6,248 data points for training and 692 for testing. Additionally, three datasets, namely Rest14 [27], Rest15 [25], and Rest16 [26] contain customer comments and restaurant rating information. Rest14 comprises 4,728 customer comments, with 3,608 for training and 1,120 for testing. Rest15 includes 1,746 reviews, with 1,204 for training and 542 for testing. Rest16, the 2016 benchmark dataset, consists of 2,364 reviews, of which 1,748 are used for training, and the remaining 616 are for testing. To enhance dataset quality, sentences with conflicting polarity or unclear meaning were removed from the Rest15 and Rest16 datasets. Table 1 shows the statistics of the four datasets.
@@ 4.2 Experimental Setup
We initialize the word embedding layer using a pre-trained GloVe vectors [23] with a dimensionality of 300. The hidden layer of the Bi-LSTM also has a dimensionality of 300. The Bi-GCN of SentiSys contains 3 layers with parameters initialized using a uniform distribution. The optimizer is Adam [39], which has a learning rate of 0.001. We use $\mathit{L_{2}}$ regularization with a batch size of 32 for up to 100 epochs.
@@ 4.3 Evaluation Measures
We use two performance metrics, Accuracy (Acc) and Macro F1 (F1), to evaluate and validate the effectiveness of the proposed SentiSys. Acc and F1 are widely recognized in the field of multi-classification tasks to assess the predictive and classification capabilities of models [35]. Concretely, Acc quantifies the ratio of correctly classified samples and thus directly measures the accuracy of the model. F1 takes into account both accuracy and recall, offering a comprehensive perspective on the model’s performance.
@@ 4.4 Evaluation Results
Table 2 presents a comparative analysis of the Acc and F1 scores between the proposed SentiSys model and 15 other baseline models. Notably, the SentiSys model demonstrates superior performance across various metrics in the Twitter, Rest14, and Rest15 datasets. While the TNet-LF model achieved the highest Acc and F1 scores in the Twitter and Rest16 datasets, it did not exhibit the same level of effectiveness as the SentiSys model in the Rest14 and Rest15 datasets. Specifically, on the Rest14 dataset, the SentiSys model stands out, achieving significantly higher Acc and F1 scores compared to the other baseline models. In the Rest15 dataset, RACL achieved the highest Acc score, but its F1 score was relatively low at 59.85. When it comes to the Rest16 dataset, the Acc and F1 scores of TNet-LF were only marginally superior to those of the SentiSys model. This observation could be attributed to the critical role of information from both parent and child nodes, highlighting the significance of considering the entire graph structure rather than solely processing it as a directed graph. Additionally, the limited sentence grammar in the Twitter dataset may have constrained prediction results. In summary, the experimental results show that our proposed SentiSys enhances the performance of aspect-based sentiment analysis systems compared with other baseline models, especially for TNet-LF on the Rest16 dataset. Furthermore, SentiSys outperforms ASGCN on all four benchmark datasets, which highlights the impact of word dependence on enhancing the classification results.
@@ 4.5 Effect of GCN Layers
Figures 3 and 3 show the effects of different GCN layers on the Acc and F1 scores of the proposed SentiSys model on the Rest14 dataset. X and Y axes indicate the number of GCN layers and the corresponding Acc and F1 scores, respectively. The experimental results show that the Acc and F1 scores tended to increase and then decrease as the number of layers increased. The Acc and F1 scores achieved the maximum values with the number of GCN layers of 3. The scores are 0.818 and 0.731, respectively.